OpenStudy (anonymous):

What is the x-intercept of the line with the equation 1/3x+y=-15 the 1/3x is a fraction

OpenStudy (anonymous):

plz help

OpenStudy (gottennis121):

@lsherron

OpenStudy (anonymous):

Slope = -0.667/2.000 = -0.333 x-intercept = -45/1 = -45.00000 y-intercept = -45/3 = -15

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome

OpenStudy (texaschic101):

1/3x + y = -15 to find the x intercept, sub in 0 for y 1/3x + 0 = -15 1/3x = -15 x = -15/(1/3) x = -15 * 3 x = - 45.....so you x intercept is -45 or (-45,0)